A wheel with radius 0.4m is mounted on an axle and is free to rotate. Beginning with the wheel at rest, a string wrapped around the edge of the wheel, is pulled with a force of 20 N . While the string is unwinding the wheel rotates through an angle of 10 rad . If the angular speed of the wheel after the string is pulled is 4 rad/s , what is the moment of inertia of the wheel?
Answer #1

r = radius = 0.4 m

F = force = 20 N

\theta = angular displacement = 10

Wi = initial angular velocity = 0 rad/s

Wf = Final angular velocity = 4 rad/s

\alpha = angular acceleration = ?

Using the equation

Wf2 = Wi2 + 2\alpha\theta

42 = 02 + 2(10)\alpha

\alpha = 0.8

Torque is given as

Fr = I\alpha

20 x 0.4 = I (0.8)

I = 10 kgm2